OpenStudy (anonymous):

please help: if g(x)=x^3+3 h(x)=3x+2 Find (3g+3h)(-x) and explain the answer please

OpenStudy (anonymous):

\[3g+3h(-x)=3g(-3)+3h(-x)=3((-x)^3+3)+3(3(-x)+2)\]

OpenStudy (anonymous):

then some algebra to get \[3(-x^3+3)+3(-3x+2)\] \[-3x^3+9-9x+6\] \[-3x^3-9x+15\]